OpenStudy (anonymous):

Solve this trigonometric equation. 2sin^2-1=0

OpenStudy (anonymous):

Note that this is the same as \[\cos (2x)=0\]

OpenStudy (anonymous):

\[\sin(x)=\pm\frac{\sqrt{2}}{2}\] is a start

OpenStudy (anonymous):

how didyou get cos(2x)

OpenStudy (anonymous):

Therefore \[x=\pi/4+n*\pi/2, n \in \mathbb{Z}\]

OpenStudy (campbell_st):

try 2sin^2x = 1 and then sin^2x = 1/2 then go from there

OpenStudy (campbell_st):

so\[\sin(x) = \sqrt{(1/2)} or \pm1/\sqrt{?}\]

OpenStudy (campbell_st):

? = 2

OpenStudy (anonymous):

but the restriction id 0 to pi/2

OpenStudy (anonymous):